Solve the system of equations by graphing. Then classify the system as consistent or inconsistent and as dependent or independent.
5x-4y=-28
4y-5x=28
What is the solution of the system of equations?
A. A point
B. No solution
C. Infinitely many solutions
Is the system consistent or inconsistent?
Are the equations dependent or independent?

Answer by drj(1380)

Solve the system of equations by graphing. Then classify the system as consistent or inconsistent and as dependent or independent.
5x-4y=-28 Equation 1
4y-5x=28 Equation 2 or writing as -5x+4y=28 multiplying by -1 to both sides leads to an identical Equation 1. That is, Equations 1 and 2 are identical.

What is the solution of the system of equations?
A. A point
B. No solution
C. Infinitely many solutions (ANSWER)

Is the system consistent or inconsistent? consistent

Are the equations dependent or independent? dependent

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J